a compound formed by the elements a and b has a cubic structure in which a atoms are the corners of the cube and b atoms are at the face centres. derive the formula of the compound. what is the co-ordination number of a and b d) what type of forces found in i) he atoms ii) nacl

Grade:12

1 Answers

Arun
6 years ago
Dear Nilesh
 
As A atoms are present at the 8 corners of the cube, therefore number of atoms of A in the unit cell is = 1/8 x 8 = 1
 
As B atoms are present at the face centres of the C faces of the cube thus no. of atoms of B = 1/2 x 6 = 3
Ratio A:B = 1:3
So formula of compound is AB3
 
 
 
Regards
Arun (askIITians forum expert)